# WallFacer: Harnessing Multi-dimensional Ring Parallelism for Efficient Long Sequence Model Training

### v3

@@ 3.3 Communication Configuration Generator
Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions. Specifically, during forward propagation, the queries do not require rearrangement; however, the keys and values must be transmitted to their corresponding positions in the ring prior to commencing the loop. As illustrated in Figure 5, this initialization ensures that each team member holds a different shard of keys and values. Moreover, it guarantees that no two teams within the same ring possess identical keys and values.
Furthermore, Team 0 and Team 1 form what is termed a team group, where ring P2P communication occurs exclusively within this group, similarly for Team 2 and Team 3. A team group is defined as a collection of teams that participate in the same sub-rings. The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$. We give the details of the determination of initial sending targets in the case of placing teams within the same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
@@
4:  target device intra-team rank = $\textbf{r}_{t}$ $\textbf{d}_{a}$
@@
After the initialization of activations, we can set up the rings by providing the GPUs their last and next GPU within their rings, as is described in Algorithm 3
@@
3:  next team in group = $(r_{t}+1)\%$ team group size $+$ team group size $\times$×self team group rank
4:  last team in group = $(r_{t}-1)\%$ team group size $+$ team group size $\times$×self team group rank
5:  next device global rank = $r_{a}+$ next team in group ×$\times d_{a}$
6:  last device global rank = $r_{a}+$ last team in group ×$\times d_{a}$
@@
3.4 Communication Topology Scheduler
The communication topology scheduler employs a grid-search algorithm to discover the optimal parallelism configuration, tailored to the specifics of the current cluster, model, and input data.
When using Ring Attention, all GPUs form a single ring, which considerably limits the flexibility of adjusting the placement scheme. In contrast, WallFacer expands the tuning space with its unique multi-ring structure in two significant ways. First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$. Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but can highly reduce the overall communication volume.
Second, the additional dimension provided by WallFacer allows for two strategies in parallelism placement. One strategy is to keep the all-gather and reduce-scatter operations intra-node, which confines these processes within the same computing node. The other strategy is to keep the ring P2P communication intra-node, reducing latency and potentially enhancing performance. The choice between these strategies depends on the extent of overlapping achievable within the ring and the communication volume during collective communications.
Based on the above tuning space, we provide an auto-scheduler, grid-searching through all possible configurations for the one that provides the highest throughput, which can be described as:
$\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$ (8)
The scheduler requires only a few iterations to profile the performance of automatically generated configurations. The time spent on this profiling is negligible compared to the entire training process, as it only needs to be performed once for each new computing cluster used for training.
3.5 Sequence Parallelism Dataloader
To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing. For full masks, sequences are straightforwardly divided into sub-sequences of equal lengths and distributed among the devices. However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end. To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by [zilin2024ringflashattention], illustrated in Figure 6. The figure illustrates the simplest case of zigzag load-balancing. Notably, the effectiveness of this strategy improves as the number of GPUs increases. This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends. This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention[li2024distflashattn].
Figure 7: Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters. We place the performance of WallFacer with both C=2 and C=4 in the figure. The configurations are marked in the titles of the sub-figures. For instance, A100_40GB_8(1B, 512K) represents that the experiment is on machines with 8 Nvidia A100 40GB GPUs in each node, the model used has one billion parameters, and the sequence length is 512k.
3.6 WallFacer Runtime
@@
WallFacer is written in PyTorch[pytorch] and uses the PyTorch torch.autograd.function and NCCL[nccl] backend for forward and backward implementation. WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
@@
Ingetrate Flash Attention. The WallFacer attention mechanism involves multiple iterations that loop over Keys and Values, with each iteration still using traditional self-attention with corresponding Query, Key, and Value (QKV). This approach enables WallFacer to incorporate flash attention effectively, extending its capability by preserving intermediate states across iterations. Additionally, WallFacer enhances the efficiency of the forward process with the help of torch JIT to fuse kernels aside from flash attention.
@@
Overlap communication with computing. In WallFacer attention, P2P communication and self-attention computing are interleaved across iterations, each incurring considerable time. To mitigate this, WallFacer employs a double buffering technique to asynchronously execute communication and computing kernels, effectively overlapping these processes and enhancing GPU utilization.
@@
Save recomputation with checkpoints. WallFacer adopts the checkpointing strategy introduced by DistFlashAttn[li2024distflashattn], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer. This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
4 Evaluation
The computational resources we use in the experiments include a local Nvidia H100 cluster with eight nodes and three Nvidia A100 clusters, as listed in table 2. We utilize two model types of total three settings, as listed in table 3. For the DiT(Diffusion Transformer) model, we use similar configurations as those in Stable Diffusion 3[esser2024sd3]. We utilize the backbone Diffusion Transformer only, without other components like the text and image encoders. During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
Table 2: Cluster Configurations. All GPUs are connected with NVLink with computing nodes. #GPU represents local size times the number of nodes.
GPU #GPUs Memory Inter-node
Nvidia H100 8*8 80GB InfiniBand
Nvidia A100 16*2 40GB 100Gbps Ethernet
Nvidia A100 8*4 40GB 100Gbps Ethernet
Nvidia A100 16*2 40GB 100Gbps Ethernet
Table 3: Model Configurations
Model Name #Heads #Layers Hidden Dim
GPT 3B 12 16 4096
GPT 7B 32 32 4096
DiT 1B 24 24 1536
In the evaluation section, we aim to answer three major questions:
  • How much improvement in throughput can WallFacer bring? Additionally, how adaptable is WallFacer to clusters with both good and poor inter-node connections?
  • Is the additional memory cost incurred by WallFacer acceptable considering the throughput improvement it offers?
  • How does WallFacer perform in scenarios of weak and strong scaling? Specifically, does it outperform Ring Attention when scaled to handle longer inputs?
Figure 8: The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
4.1 Throughput and Adaptability
Our first experiment aims to assess the performance of WallFacer and Ring Attention across different clusters with varying environments, testing the adaptability of both methods. There are several factors influencing the efficiency of ring-style attention computation:
Theoretical Computation-Communication Volume Ratio: Primarily determined by the sequence length used during training. Attention computation exhibits a computational complexity of $O(N^{2}\cdot H)$, whereas P2P communication complexity is $O(N\cdot H)$. Thus, the model configuration does not impact this ratio; only the sequence length does. A larger $N$ increases the computation-communication ratio, facilitating easier overlap of communication with computation. We evaluate varying sequence lengths to explore performance under different ratios.
Compute Capability and Connectivity of GPUs: The computing overhead, given a specific volume, affects the computation-communication overhead ratio. Higher compute capabilities make overlapping more challenging. We utilize two sets of GPUs in this evaluation: Nvidia A100 40GB and Nvidia H100 80GB, with the latter offering significantly higher theoretical tflops on bf16 computations. Connectivity is considered in two parts: intra-node and inter-node. Our clusters are equipped with NVLink, ensuring robust intra-node communication. For inter-node communication, while InfiniBand offers high bandwidth at a higher commercial cost, some clusters utilize Ethernet for connectivity. Specifically, our H100 nodes leverage InfiniBand with eight adapters per node for superior inter-node bandwidth, whereas the Google Cloud servers use Ethernet. The diversity in node configurations (8-GPU and 16-GPU nodes) allows us to assess adaptability across different topologies.
This evaluation not only highlights the inherent differences between the schemes but also tests their flexibility in various hardware settings.
The results of our evaluation are illustrated in Figure 7. We measure throughput in thousands of tokens per second. To better demonstrate how to select the optimal configuration of WallFacer under each condition, we included two configurations, Wall-2 and Wall-4, in the figure. We omit configurations with lower performance for clarity. As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 114.33%, 41.4%, 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%. This advantage is primarily due to the additional parallel dimension that WallFacer introduces. Unlike Ring Attention, which requires inter-node P2P communication in each iteration, WallFacer’s P2P communication is mostly confined intra-node, except for initial data transfers. This experiment clearly demonstrates WallFacer’s superior performance across various environments.
Another observation is that the optimal configuration for WallFacer may vary depending on the environment, reflecting differences in the computation-communication ratio and the trade-offs between collective and P2P communication. For example, the best $C$ value for A100_16 is 2, while for A100_8, it is 4. This variation can be attributed to the topological differences: the 16-GPU-node cluster benefits less from reduced P2P communication volume, making a smaller $C$ preferable to decrease collective communication volume while maintaining P2P communication at an acceptable level. This finding underscores the importance of the Communication Topology Scheduler we provide, which helps users avoid the complex process of manually analyzing these factors.
In summary, WallFacer demonstrates superior efficiency and adaptability in all six cases, thanks to its flexible parallelism scheme.
@@ 4.1 Throughput and Adaptability
Table 4: Supported Sequence Length of Ring Attention and WallFacer on one Nvidia A100 80GB GPU.
@@
Model Size Length Ring Attention WallFacer
@@
4.2 Memory Consumption
Before evaluating the memory consumption of WallFacer, we first compared the supported sequence lengths of WallFacer with those reported in the RingAttention paper [liu2023ring]. As shown in Table 4, despite the slightly higher memory requirements of WallFacer, it still supports sequence lengths commonly used in training tasks.
Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation. To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch [pytorch] during our experiments. It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency. To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer, except the DiT 512k Wall-4 setting. However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable. We observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
Moreover, in scenarios involving larger models, the relative increase in memory consumption due to QKV duplication diminishes. This reduction is explained by equation 7, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant. This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B and 1B model.
4.3 Strong and Weak Scaling
In the scaling tests we carry out experiments for both strong and weak scaling. Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up. So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64 for the GPT model and from 8 to 32 for the DiT model. As id depicted in Figure LABEL:Fig.strong, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs. When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention for the GPT model and 114.33% for the DiT model on 32 A100 GPUs. This can also explained by the computation-communication ratio. When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation. This is exactly the main advantage of WallFacer, reduction of total P2P communication volume. So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
In our weak scaling experiments using DiT models, we scale the sequence length from 128k to 512k while proportionally increasing the number of GPUs from 8 to 32. Assuming a scaling factor of $k$, both the sequence length and the number of GPUs increase by $k$, resulting in a total computation volume that increases by approximately $k^{2}$. Since the computational resources also increase by $k$, the theoretical overall throughput (tokens per second) should remain constant during this weak scaling setting. As shown in Figure LABEL:Fig.weak, both WallFacer and Ring Attention exhibit this expected behavior, with throughput remaining relatively constant. However, WallFacer consistently achieves higher throughput across different numbers of GPUs. This is because, although the local computation-to-communication ratio stays the same in weak scaling, Ring Attention requires a higher proportion of inter-node communication as the number of GPUs increases, whereas WallFacer keeps most communication local.
In summary, WallFacer shows better scalability in both strong and weak scaling experiments, making it a better choice for large-scale Transformer model training.
5 Discussion
5.1 FlashAttention3 and Hopper GPUs
@@
In addition to the original FlashAttention [dao2022flashattention] used in our experiments, FlashAttention3 [shah2024flashattention3fastaccurateattention] has been introduced, specifically designed for Hopper and newer Nvidia GPUs. For FP16 precision, which is utilized in this paper, FlashAttention3 achieves a 1.5-2.0x speedup on Hopper GPUs. As discussed in Section 2, reducing attention computation overhead results in more P2P communication not being overlapped, further emphasizing the need to reduce communication volume. With the increasing adoption of Hopper GPUs, the significance of the WallFacer system will also grow.
5.2 WallFacer and Other Parallelisms
@@
Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with WallFacer with minimal effort. However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like Ulysses. Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making WallFacer orthogonal to it.
@@
Other Sequence Parallelism WallFacer is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed Ulysses [jacobs2023Ulysses]. While Ulysses distributes attention heads across different devices, WallFacer can independently partition activations along the sequence length dimension. In future work, we can explore combining WallFacer with DeepSpeed Ulysses to expand the communication scheduling space, harnessing the scalability of WallFacer alongside the efficiency of Ulysses.
@@
In summary, WallFacer can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
6 Related Work
@@
Attention Optimization. Traditional full attention mechanisms necessitate $O(n^{2})$ memory for storing the outputs of $QK^{T}$, leading to significant computational and memory demands. To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements. Memory-efficient attention[rabe2022selfattention] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory. Flash Attention further minimizes I/O overhead and enhances overall efficiency. Additionally, optimization methods specifically tailored for inference, such as PagedAttention[kwon2023efficient], are also being developed to improve the efficiency of attention computations. In this work, we utilize Flash Attention within each iteration to reduce the computation overhead.
@@
Long-Sequence Training Techniques. Sequence Parallelism[Li2021SequencePL] was initially introduced to enhance the efficiency of parallel long-sequence training. Ring Attention[liu2023ring] improved communication efficiency through memory-efficient methods[rabe2022selfattention], supporting near-infinite sequence lengths. DeepSpeed Ulysses[jacobs2023Ulysses] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads. Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[li2024distflashattn] features a load-balance scheme and a novel gradient checkpoint method. Our work builds on these innovations, introducing a system that supports large-scale training with an efficient communication scheme.
@@
Techniques for Distributed Model Training. Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[micikevicius2018mixed] and the ZeRO series[rajbhandari2020zero]. In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently. 2) Hybrid Parallelism: Frameworks like Megatron[megatron] and Colossal AI[bian2021colossal] integrate multiple forms of parallelism. There are various existing Parallelism techniques like Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] and Tensor Parallelism[shoeybi2019megatron], which can be combined with WallFacer Parallelism to facilitate large-scale training. We are also considering the integration of additional frameworks such as [chang2024centauri] to enhance overlapping capabilities in future implementations.
7 Conclusion
WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme. Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions for both Computer Vision(CV) and Natral Language Processing(NLP) models. In an era increasingly demanding longer contexts for both NLP and CV, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.
[…]

### v4

@@ A.2 Details in the training process
Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions. Specifically, during forward propagation, the queries do not require rearrangement; however, the keys and values must be transmitted to their corresponding positions in the ring prior to commencing the loop. As illustrated in Figure 6 and algorithn, this initialization ensures that each team member holds a different shard of keys and values. Moreover, it guarantees that no two teams within the same ring possess identical keys and values. Initially, for the setup stage, it is essential to establish the sub-rings by rearranging the activation positions. Specifically, during forward propagation, the queries do not require rearrangement; however, the keys and values must be transmitted to their corresponding positions in the ring prior to commencing the loop. As illustrated in Figure 6 and algorithm, this initialization ensures that each team member holds a different shard of keys and values. Moreover, it guarantees that no two teams within the same ring possess identical keys and values.
@@
4: target device intra-team rank = $\textbf{r}_{t}$ % $\textbf{d}_{a}$
@@
After the initialization of activations, we can set up the rings by providing the GPUs their last and next GPU within their rings, as is described in Algorithm 3.
@@
3: next team in group = $(r_{t}+1)\%$ team group $size ×\timesself team group$ rank
4: last team in group = $(r_{t}-1)\%$ team group $size ×\timesself team group$ rank
5: next device global rank = $r_{a}+$ next team in group $\times d_{a}$
6: last device global rank = $r_{a}+$ last team in group $\times d_{a}$
@@
A.3 StarTrail Runtime
@@
StarTrail is written in PyTorch[29] and uses the PyTorch torch.autograd.function and NCCL[28] backend for forward and backward implementation. StarTrail also employs multiple techniques during runtime to improve its overall training efficiency.
@@
Ingetrate Flash Attention. The StarTrail attention mechanism involves multiple iterations that loop over Keys and Values, with each iteration still using traditional self-attention with corresponding Query, Key, and Value (QKV). This approach enables StarTrail to incorporate flash attention effectively, extending its capability by preserving intermediate states across iterations. Additionally, StarTrail enhances the efficiency of the forward process with the help of torch JIT to fuse kernels aside from flash attention.
@@
Overlap communication with computing. In StarTrail attention, P2P communication and self-attention computing are interleaved across iterations, each incurring considerable time. To mitigate this, StarTrail employs a double buffering technique to asynchronously execute communication and computing kernels, effectively overlapping these processes and enhancing GPU utilization.
@@
Save recomputation with checkpoints. StarTrail adopts the checkpointing strategy introduced by DistFlashAttn[20], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer. This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
Appendix B Additional Experiment
To comprehensively evaluate the memory consumption of StarTrail and Ring Attention, we compared the maximum supported sequence lengths of StarTrail with those reported in the Ring Attention paper [24]. As shown in Table 2, although StarTrail requires slightly more memory, it still supports sequence lengths commonly used in training tasks.
@@ Appendix B Additional Experiment
Table 2: Supported Sequence Length of Ring Attention and StarTrail on one Nvidia A100 80GB GPU.
@@
Model Size Length Ring Attention StarTrail
@@
B.1 Discussion
B.2 Larger Batch Sizes and Models
We utilized small batches and model sizes in our experiments because these choices do not affect the underlying improvements in communication efficiency and computation-to-communication ratio that StarTrail provides. For larger batch sizes, both communication and computation scale proportionally, leaving the overlapping ability unchanged. Similarly, while larger models involve more layers or larger hidden sizes, the key attention computations and corresponding ratios remain unaffected. Hence, our conclusions naturally extend to scenarios with larger batches and models.
B.2.1 FlashAttention3 and Hopper GPUs
@@
In addition to the original FlashAttention [8] used in our experiments, FlashAttention3 [35] has been introduced, specifically designed for Hopper and newer Nvidia GPUs. For FP16 precision, which is utilized in this paper, FlashAttention3 achieves a 1.5-2.0x speedup on Hopper GPUs. As discussed in Section 4, reducing attention computation overhead results in more P2P communication not being overlapped, further emphasizing the need to reduce communication volume. With the increasing adoption of Hopper GPUs, the significance of the StarTrail system will also grow.
B.2.2 StarTrail and Other Parallelisms
@@
Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with StarTrail with minimal effort. However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like DeepSpeed-Ulysses. Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making StarTrail orthogonal to it.
@@
Other Sequence Parallelism StarTrail is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed-Ulysses [14]. While DeepSpeed-Ulysses distributes attention heads across different devices, StarTrail can independently partition activations along the sequence length dimension. In future work, we can explore combining StarTrail with DeepSpeed-Ulysses to expand the communication scheduling space, harnessing the scalability of StarTrail alongside the efficiency of DeepSpeed-Ulysses.
@@
In summary, StarTrail can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
B.3 Other Related Works
@@
Attention Optimization. Traditional full attention mechanisms necessitate $O(n^{2})$ memory for storing the outputs of $QK^{T}$, leading to significant computational and memory demands. To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements. Memory-efficient attention[32] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory. FlashAttention further minimizes I/O overhead and enhances overall efficiency. Additionally, optimization methods specifically tailored for inference, such as PagedAttention[19], are also being developed to improve the efficiency of attention computations. In this work, we utilize FlashAttention within each iteration to reduce the computation overhead.
@@
Long-Sequence Training Techniques. Sequence Parallelism[21] was initially introduced to enhance the efficiency of parallel long-sequence training. Ring Attention[24] improved communication efficiency through memory-efficient methods[32], supporting near-infinite sequence lengths. DeepSpeed-Ulysses[14] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads. Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[20] features a load-balance scheme and a novel gradient checkpoint method. Our work builds on these innovations, introducing a system that supports large-scale training with an efficient communication scheme.
@@
Techniques for Distributed Model Training. Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[26] and the ZeRO series[34]. In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently. 2) Hybrid Parallelism: Frameworks like Megatron[27] and Colossal AI[4] integrate multiple forms of parallelism. There are various existing Parallelism techniques like Pipeline Parallelism[13, 10, 23, 25] and Tensor Parallelism[37], which can be combined with StarTrail Parallelism to facilitate large-scale training. We are also considering the integration of additional frameworks such as [6] to enhance overlapping capabilities in future implementations.